a canoeist can travel 10 mph in calm water. the amount of time it takes to travel 60 miles with the current is the same amount of time it takes to travel 40 miles against the current. find the rate of the current.
Answer by checkley71(8403)

SET UP A FRACTIONAL RELATIONSHIP
10/60=X/40
60X=400
X=400/60
X=6/6666667 OR 6 2/3 MPH IS THE CURRENT